Give the gas you are assigned here And give your
Answer #1

Since, it is an ideal gas, it follows the equation, PV = nRT, P = 2.5 atm, n = 3, T = 273.15, R = 0.0821 L atm/mol/K.

Putting these values in the equation, V = nRT/P = 26.91 L

van Der Waal's eq = (P + an^2/V^2)(V-nb) = nRT , a = 0.034, b = 0.023

Putting these values in eqn, we find V = 26.977 L